A force of 600 N stretches a certain spring a distance of 0.300 m .

Part A)

What is the potential energy of the spring when it is stretched a distance of 0.300 m ?

Part B)

What is its potential energy when it is compressed a distance of 6.00 cm ?
Answer #1

We know that Magnitude of force F=KX

K=F/X=600/0.3

K=2000N/m

PART A)

We know that potential energy of the spring

U=KX2 /2

here K=2000 N/m , X=0.300 m

U=2000*(0.3)2 /2

U=90 J Answer.

part B)

here X=6.00 cm=0.06 m

U=2000*(0.06)2 /2

U=3.60 J Answer